# 19B Yuquan Road, Shijingshan District, Beijing 100049, China

### v1

@@ 1 Introduction
2 Maximizing CMI: the Multi-entanglement phase transition theorem
@@ 2.1.1 $E$ being one interval
which does not depend on the value of $X(A:B)$. Note that the condition of a cross ratio being $1$ is just the critical point for the entanglement phase transition between the connected and disconnected phase of the corresponding MI as can be seen from equation (2.4). Therefore, in order to reach the upper bound of CMI for the conditional region $E$ being one single interval, one should adjust the interval $E$ to reach the critical points of both the entanglement phase transitions with $A$ and the one with $B$ simultaneously.
@@ 3.1 MPT diagram
  • $m=2$: phase transitions $A/E_{1}$, $B/E_{2}$, $AE_{1}/E_{2}$, and $BE_{2}/E_{1}$ occur simultaneously.
@@
The third constraint, i.e., the disconnectivity condition, which states the vanishing of mutual information $I(A:E)$ and $I(B:E)$ when CMI reaches its upper bound, is a very powerful constraint. We will revisit it in Section 4 as a core statement. For now, we can intuitively understand this statement as follows. In Section 2.2, when we try to prove the multi-entanglement phase transition rule, we only need to analyze the phase transition critical conditions for RT surfaces of regions that appear in formula 1, e.g., $AE$, $BE$, $E$, and $ABE$. The entanglement wedge of region $AE$ will not be affected by the phase transition of $A/E_{i}$ [^4] Note that we are saying that the connectivity of $A$ and $E_{i}$ in the entanglement wedge of $AE_{i}$ does not matter, while the connectivity of $A$ and $E_{i}$ in the entanglement wedge of the whole $AE$ still matters.; only when the whole entanglement wedge $AE$ reaches its phase transition critical point could CMI be affected. In other words, MPT diagrams with conditions like $A/E_{i}$ and $A/E_{j}$ will not be the diagrams with maximum CMI. Only conditions like $AE_{i_{1}}E_{i_{2}}.../E_{i_{m}}$ are valid. In Figure 9, only diagrams on the first row satisfy this disconnectivity condition. In practice, one can indeed find that $I_{1}$ and $I_{2}$, which are much larger than $I_{3}$ and $I_{4}$, are the only two rightful candidates.
@@ 4 The upper bound of CMI: a more universal method to determine E
In the previous section, we analyze the divergence behaviour of $I_{3}$ and come to the conclusion that nearly all UV degrees of freedom of $A$ and $B$ are participating in the tripartite entanglement. However, as we cannot determine the exact relationship between the number of intervals $m$ and the UV cutoff $\epsilon$, the $O(1)$ term in formula (30) is still unknown. In this section, we are trying to analyze whether all the IR degrees of freedom are participating in tripartite entanglement as well. We are going to investigate more complicated cases, such as the two sided black hole and higher-dimensional cases. In higher dimensions, the MPT theorem does not hold true any more; however, the disconnectivity condition is still valid in a weaker form. We first introduce this weaker form of the disconnectivity condition and prove it. Then we utilize this condition to give a direct calculation of the upper bound of CMI in various general cases. At last, we give a general and concise formula for the upper bound of CMI.
@@ 4.1 Disconnectivity condition
The third step: we continuously enlarge the gap inside $E$. One can find that $S_{E_{i1}}$ and $S_{E_{i2}}$ both decrease, while the entropy of the gap region in-between $S_{gap}$ increases. These three quantities are the only ones that affect the value of CMI because the outer boundary of $E_{i}$ remains unchanged. The sign of them inside formula (1) is drawn as the different colors of those RT surfaces in Figure 17; red represents positive and blue represents negative. It is easy to observe that when we enlarge the gap, CMI increases until another phase transition occurs. In principle, the phase transition could occur on one of $E_{i1}$ or $E_{i2}$, i.e.the entanglement wedge of $AE$ [^8] Here, we assume that $E_{i1}$ and $E_{i2}$ become disconnected in the entanglement wedge of $AE$ before $E_{i1}$ and $E_{i2}$ become disconnected in the entanglement wedge of $BE$, without loss of generality. suddenly becomes disconnected between $A$ and $E_{i1(2)}$ while remaining the connectivity between $A$ and $E_{i2(1)}$. However, we can make the phase transition occur on $E_{i1}$ and $E_{i2}$ simultaneously by adjusting the middle point of this gap [^9] If the middle point of the gap region between $E_{i1}$ and $E_{i2}$ is far left inside the original region $E$, $E_{i1}$ disconnects with $A$ first; otherwise, if the middle point is far right inside the original region $E$, $E_{i2}$ disconnects with $A$ first. Therefore, there must exist a fine-tuned middle point which makes the phase transition occur on $E_{i1}$ and $E_{i2}$ simultaneously.. After the phase transition, $E_{i1}$ and $E_{i2}$ disconnect with $A$ in the entanglement wedge of $AE$. Remember that when we enlarge the gap, the CMI increases, so we prove that: for a configuration whose intervals of $E$ connect with both $A$ and $B$ inside the entanglement wedge of $AE$ and $BE$ respectively, one can always find another configuration whose intervals connect to at most one of $A$ and $B$ inside the entanglement wedge of $AE$ and $BE$, with larger CMI.
@@ 4.3 Entanglement of state-constrained purification
We would like to impart the area of the minimal surface $S_{D}$ a physical meaning in the sense of quantum information theory. The definition of the RT surface of $D$ is very similar to the definition of the entanglement wedge cross section (EWCS). In Takayanagi:2017knl , EWCS between $A$ and $B$ is defined as the minimal surface that divides the entanglement wedge of $AB$ into two parts, with its area being conjectured as the entanglement of purification (EoP) between $A$ and $B$. When the entanglement wedge of $AB$ is disconnected, no surface is needed to divide it into two parts. The area of EWCS vanishes in this case Jain:2022csf , as does the EoP when $I(A:B)=0$ to leading order in $\frac{1}{G_{N}}$.
@@ 5.1 Overview
Can we draw MPT diagrams as we did in Section 3 to analyze $I_{4}$? Unfortunately, the answer is no, because given a finite $m$, the configuration with maximum $I_{4}$ is not unique. Unlike the case we presented in Section 2.1, the maximum value of $I_{4}$ is not a peak but a plateau on the mountain, which means $I_{4}$ remains maximum in a set of configurations rather than in a unique MPT diagram. The MPT theorem does not hold true in the $I_{4}$ case. Therefore we have to employ the disconnectivity condition to find the upper bound configuration for $I_{4}$.
@@ Acknowledgement
We would like to thank Bartłomiej Czech and Joydeep for his valuable discussions. This work was supported by Project 12347183, 12035016 and 12275275 supported by the National Natural Science Foundation of China. It is also supported by Beijing Natural Science Foundation under Grant No. 1222031.

### v2

@@ 1 Introduction
It is worth noting that the multi-partite entanglement we investigate is only one type of the multipartite entanglement structure that could be revealed by the entanglement entropy of the subsystems. However, even if two multipartite quantum systems have all subsystems with the same entanglement entropy, those two systems could still not be local unitary equivalent, or could not transform into each other by local operations and classical communication (LOCC) bengtsson2016brief . To distinguish such delicate difference between those two quantum systems, measures which are beyond the combination of entanglement entropy should be introduced, such as reflected entropy Dutta:2019gen , negativity Calabrese:2012ew ; Kusuki:2019zsp , or more general multipartite measures Gadde:2022cqi ; Penington:2022dhr ; Yuan:2024yfg , etc. Those topics are beyond the concept of this paper.
@@
2 Maximizing CMI: the multi-entanglement phase transition rule
@@ 2.1.1 $E$ being one interval
which does not depend on the value of $X(A:B)$. Note that the condition of a cross ratio being $1$ is just the critical point for the entanglement phase transition between the connected and disconnected phase of the corresponding MI as can be seen from equation (5). Therefore, in order to reach the upper bound of CMI for the conditional region $E$ being one single interval, one should adjust the interval $E$ to reach the critical points of both the entanglement phase transitions with $A$ and the one with $B$ simultaneously. Note that it is always possible to satisfy both of these conditions, because there are two degrees of freedom of region $E$ (the two coordinates of the end points of region $E$), and those two entanglement phase transition conditions are two equations that uniquely determine them. Intuitively, one can choose the middle point of region $E$ inside the gap, then enlarge $E$ while preserving its middle point. If the middle point is near $A$ and far from $B$, $EW(AE)$ will first connect during the enlarging process and vise versa. As a result, because of the intermediate value theorem, there must exist a fine-tuned middle point of $E$ that makes $EW(AE)$ and $EW(BE)$ connect simultaneously, which is the $E$ that maximizes the CMI.
@@ 3.1 MPT diagram
  • $m=2$: phase transitions $A/E_{1}$, $B/E_{2}$, $AE_{1}/E_{2}$ [^4] In Section 2, we write the entanglement phase transition condition as $AD_{1}E_{1}/E_{2}$, while we omit the gap region $D_{1}$ in between region $A$ and $E_{1}$ here. Technically, this entanglement phase transition condition only constrains the four coordinates of the endpoints $\partial AD_{1}E_{1}$ and $\partial E_{2}$, so as long as the entanglement wedge of $AE_{1}$ is connected, the exact position of $\partial D_{1}$ is irrelevant in condition $AD_{1}E_{1}/E_{2}$. Therefore, the two conditions $AD_{1}E_{1}/E_{2}$ and $AE_{1}/E_{2}$ are in fact equivalent and as a result, for convenience, we can simply write $AE_{1}/E_{2}$ instead., and $BE_{2}/E_{1}$ occur simultaneously.
@@
The third constraint, i.e., the disconnectivity condition, which states the vanishing of mutual information $I(A:E)$ and $I(B:E)$ when CMI reaches its upper bound, is a very powerful constraint. We will revisit it in Section 4 as a core statement. For now, we can intuitively understand this statement as follows. In Section 2.2, when we try to prove the multi-entanglement phase transition rule, we only need to analyze the phase transition critical conditions for RT surfaces of regions that appear in formula 1, e.g., $AE$, $BE$, $E$, and $ABE$. The entanglement wedge of region $AE$ will not be affected by the phase transition of $A/E_{i}$; [^5] Note that we are saying that the connectivity of $A$ and $E_{i}$ in the entanglement wedge of $AE_{i}$ does not matter, while the connectivity of $A$ and $E_{i}$ in the entanglement wedge of the whole $AE$ still matters. only when the whole entanglement wedge $AE$ reaches its phase transition critical point could CMI be affected. In other words, MPT diagrams with conditions like $A/E_{i}$ and $A/E_{j}$ will not be the diagrams with maximum CMI. Only conditions like $AE_{i_{1}}E_{i_{2}}.../E_{i_{m}}$ are valid. In Figure 9, only diagrams on the first row satisfy this disconnectivity condition. In practice, one can indeed find that $I_{1}$ and $I_{2}$, which are much larger than $I_{3}$ and $I_{4}$, are the only two rightful candidates.
@@ 4 The upper bound of CMI: a more universal method to determine E
In the previous section, we analyze the divergence behaviour of $I_{3}$ and come to the conclusion that nearly all UV degrees of freedom of $A$ and $B$ are participating in the tripartite entanglement. However, as we cannot determine the exact relationship between the number of intervals $m$ and the UV cutoff $\epsilon$, the $O(1)$ term in formula (30) is still unknown. In this section, we are trying to analyze whether all the IR degrees of freedom are participating in tripartite entanglement as well. We are going to investigate more complicated cases, such as the two sided black hole and higher-dimensional cases. In higher dimensions, $\partial E$ has infinitely many degrees of freedom, so any finite number of entanglement phase transition conditions cannot fully determine the shape of region $E$. In this sense, the multi-entanglement phase transition rule which we state in Section 2.2 is not useful any more; however, the disconnectivity condition is still valid in a weaker form. We first introduce this weaker form of the disconnectivity condition and prove it. Then we utilize this condition to give a direct calculation of the upper bound of CMI in various general cases. At last, we give a general and concise formula for the upper bound of CMI.
@@ 4.1 Disconnectivity condition
The third step: we continuously enlarge the gap inside $E$. One can find that $S_{E_{i1}}$ and $S_{E_{i2}}$ both decrease, while the entropy of the gap region in-between $S_{gap}$ increases. These three quantities are the only ones that affect the value of CMI because the outer boundary of $E_{i}$ remains unchanged. The sign of them inside formula (1) is drawn as the different colors of those RT surfaces in Figure 17; red represents positive and blue represents negative. It is easy to observe that when we enlarge the gap, CMI increases until another phase transition occurs. In principle, the phase transition could occur on one of $E_{i1}$ or $E_{i2}$, i.e.the entanglement wedge of $AE$, [^9] Here, we assume that $E_{i1}$ and $E_{i2}$ become disconnected in the entanglement wedge of $AE$ before $E_{i1}$ and $E_{i2}$ become disconnected in the entanglement wedge of $BE$, without loss of generality. suddenly becomes disconnected between $A$ and $E_{i1(2)}$ while remaining the connectivity between $A$ and $E_{i2(1)}$. However, we can make the phase transition occur on $E_{i1}$ and $E_{i2}$ simultaneously by adjusting the middle point of this gap [^10] If the middle point of the gap region between $E_{i1}$ and $E_{i2}$ is far left inside the original region $E$, $E_{i1}$ disconnects with $A$ first; otherwise, if the middle point is far right inside the original region $E$, $E_{i2}$ disconnects with $A$ first. Therefore, there must exist a fine-tuned middle point which makes the phase transition occur on $E_{i1}$ and $E_{i2}$ simultaneously.. After the phase transition, $E_{i1}$ and $E_{i2}$ disconnect with $A$ in the entanglement wedge of $AE$. Remember that when we enlarge the gap, the CMI increases, so we prove that: for a configuration whose intervals of $E$ connect with both $A$ and $B$ inside the entanglement wedge of $AE$ and $BE$ respectively, one can always find another configuration whose intervals connect to at most one of $A$ and $B$ inside the entanglement wedge of $AE$ and $BE$, with larger CMI.
@@ 4.3 Entanglement of state-constrained purification
We would like to impart the area of the minimal surface $S_{D}$ a physical meaning in the sense of quantum information theory. The definition of the RT surface of $D$ is very similar to the definition of the entanglement wedge cross section (EWCS). In Takayanagi:2017knl ; Mori:2024gwe , EWCS between $A$ and $B$ is defined as the minimal surface that divides the entanglement wedge of $AB$ into two parts, with its area being conjectured as the entanglement of purification (EoP) between $A$ and $B$. When the entanglement wedge of $AB$ is disconnected, no surface is needed to divide it into two parts. The area of EWCS vanishes in this case Jain:2022csf , as does the EoP when $I(A:B)=0$ to leading order in $\frac{1}{G_{N}}$.
@@ 5.1 Overview
Can we draw MPT diagrams as we did in Section 3 to analyze $I_{4}$? Unfortunately, the answer is no, because given a finite $m$, the configuration with maximum $I_{4}$ is not unique. Unlike the case we presented in Section 2.1, the maximum value of $I_{4}$ is not a peak but a plateau on the mountain, which means $I_{4}$ remains maximum in a set of configurations rather than in a unique MPT diagram. The multi-entanglement phase transition rule does not hold true in the $I_{4}$ case. Therefore we have to employ the disconnectivity condition to find the upper bound configuration for $I_{4}$.
@@ Acknowledgement
We would like to thank Bartłomiej Czech and Joydeep Naskar for theirhis valuable discussions. This work was supported by Project 12347183, 12035016 and 12275275 supported by the National Natural Science Foundation of China.